# ACTIVE: Continuous Similarity Search for Vessel Trajectories

### v1

@@ Abstract.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX Tiantian Liu and Hengyu Liu have equal contributions.
@@
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Automatic Identification System (AIS) data, which provides real-time information on vessel positions, speeds, and directions, offers a rich resource for analyzing vessel movements. With the growing availability of AIS data, understanding vessel trajectories has become essential for navigation safety (Li et al., 2024b) and environmental monitoring (Jin et al., 2021). Trajectory similarity search (Hu et al., 2023)—identifying vessels following similar paths—plays a key role in applications such as collision avoidance (Li et al., 2022), route optimization (Xu et al., 2024), and movement pattern analysis (Hu et al., 2024).
@@ 2.1. Trajectory
A segment $s_{i}=\langle p_{i}^{j^{\prime}},\dots,p_{i}^{j}\rangle$ of a trajectory $T_{i}$ is a continuous sub-sequence of $T_{i}$ defined by an index range $[j^{\prime},j]$, where $1\leq j^{\prime}\leq j\leq n$.
@@ 2.2. Problem Definition
Given a moving object $o$, a set of historical trajectories $\mathcal{T}$, a range $r$, an integer $k$, and a similarity measure $\mathcal{D}(\cdot,\cdot)$, CSTS$(o,\mathcal{T},r,k)$ continuously identifies the $k$ most similar trajectories $\Theta_{t}$ for $o$ at each timestamp $t$, where $o_{t}=(p_{o}^{c},T_{o})$, satisfying $\forall T\in\Theta_{t}\,(T\in\mathcal{T}_{r})\wedge\forall T^{\prime}\notin\Theta_{t}\,(\mathcal{D}(o,T^{\prime})\geq\mathcal{D}(o,T))$.
@@ 2.3. Framework Overview
Segment-based Vessel Trajectory Index module. Historical trajectories are partitioned into smaller segments using an MBR-optimization method and organized in an R-tree based index. This index captures connection information, enabling efficient reconstruction of full trajectories when needed and supports spatially optimized search.
Similarity Computation module. This module computes the real-time distance between a moving object and historical trajectories using the novel similarity measure OTRD, which consists of the Historical Trajectory Distance (HTD) and the Target-Trajectory Distance (TTD). The OTRD measure prioritizes the future trends of vessel movements and the target destination while also considering the relevance of trajectory segments and the weighted significance of individual points.
@@ 3.2. Segment Partitioning
Algorithm 1 outlines the segment partitioning process in detail. Initially, a segment set $\mathcal{S}$ is initialized (line 1). For each trajectory $T_{i}$, the algorithm identifies the optimal segmentation (lines 2–16). Specifically, a list is created for $T_{i}$’s segment set, where the $j$th item represents the current optimal segment set up to point $p^{j}$ (line 3). Additionally, another list is initialized to store the total MBR area, where the $j$th item represents the minimum total MBR area up to point $p^{j}$ (line 4). For each point $p^{j}$ in $T_{i}$, the minimum total MBR area up to $p^{j}$ is computed (lines 5–15). If $j$ is less than $l_{\textit{min}}$, the area is set to 00; otherwise, it is initialized to $\infty$ (lines 6–7). For each possible starting point $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$, the MBR area of the segment $\langle p^{j-l+1},...,p^{j}\rangle$ is computed (lines 8–10). If the total MBR area is smaller than the current value, both the area and the segment set are updated (lines 11–15). Once all points in $T_{i}$ have been processed, the final segment set for $T_{i}$ is added to the overall segment set $\mathcal{S}$ (line 16). Ultimately, the algorithm returns the overall segment set $\mathcal{S}$ (line 17).
2:Output: A segment set $\mathcal{S}$.
3:Initialize a segment set $\mathcal{S}$leftarrow $\emptyset$
4:for $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$ in $\mathcal{T}$ do
5:    Initialize a list for $T_{i}$’s segment set $S_{i}$leftarrow $\emptyset$
6:    Initialize a list for the total MBR area $F$leftarrow $\emptyset$
7:    for $j$ in $[1,n]$ do
8:         if $j\leq l_{\textit{min}}$ then $F[j]$leftarrow 00
9:         else $F[j]$leftarrow $\infty$
10:         for $l$ in $[l_{\textit{min}},l_{\textit{max}}]$ do
11:             if $j-l\geq 0$ then
12:                 $m_{l}$leftarrow $\mathcal{M}_{area}$($p_{i}^{j-l+1}$, $p_{i}^{j}$)
13:                 if $j-l$ is 00 then $m$leftarrow $m_{l}$
14:                 else $m$leftarrow $F[j-l]+m_{l}$
15:                 if $F[j]>m$ then
16:                     $F[j]$leftarrow $m$
17:                     $S_{i}[j]$leftarrow $S_{i}[j-l]$cup $\langle p_{i}^{j-l+1},...,p_{i}^{j}\rangle$
18:    $\mathcal{S}\leftarrow\mathcal{S}\cup S_{i}[n]$
19:return $\mathcal{S}$
@@ 4.2. Object-Trajectory Real-time Distance
(3) $\mathcal{D}_{H}(T_{o},T^{\prime}_{i})=\begin{cases}d(p_{o}^{*},p_{i}^{*}),\mbox{ if }T_{o}=\{p_{o}^{*}\}\mbox{ and }T^{\prime}_{i}=\{p_{i}^{*}\}\\ \infty,\mbox{ if }T_{o}=\emptyset\mbox{ or }T^{\prime}_{i}=\emptyset\\ \textit{max}\{d(p_{o}^{*},p_{i}^{*}),\textit{min}\{\theta\cdot\mathcal{D}_{H}(T_{o}^{*},{T^{\prime}_{i}}^{*}),\\ \mbox{ }\mbox{ }\mbox{ }\mbox{ }\mbox{ }\mbox{ }\mathcal{D}_{H}(T_{o},{T^{\prime}_{i}}^{*})\}\},\mbox{ otherwise}\end{cases}$
@@ 5.1. Speed-up Strategies
(12) $T^{\prime}_{i}(g)=\begin{cases}\langle p_{i}^{j},...,p_{i}^{x-g},p_{i}^{x}\rangle&\mbox{if }\gamma=0\\ \langle p_{i}^{j},...,p_{i}^{x-\gamma-g},p_{i}^{x-\gamma},p_{i}^{x}\rangle&\mbox{if }\gamma>0\mbox{ and }x>x^{\prime}\\ \langle p_{i}^{j},...,p_{i}^{x+\gamma-g},p_{i}^{x}\rangle&\mbox{if }\gamma>0\mbox{ and }x<x^{\prime}\end{cases}$
@@ 5.2. OTRD Computation
Algorithm 2 shows the details of the similarity computation. Specifically, we first identify the nearest segment $s_{i}^{z}$ and the nearest point $p_{i}^{x}$ on the historical trajectory $T_{i}$ to the moving object’s current position $p_{o}^{c}$ (lines 1–2). We initialize a boolean flag isValid to check if the pre-computed distance in the intermediate buffer $B_{I}$ is still valid (line 3). A variable $p_{\textit{max}}$ is also initialized to represent the ID of the point $p_{o}^{y}$ in $T_{o}$, where $\theta^{c-y}*\textit{min}\{d(p_{o}^{y},p_{i})|p_{i}\in T^{\prime}_{i}\}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ (line 3). If the pivotal point ID of $T_{i}$ increases and the distance between $p_{o}^{y^{\prime}}$ and the incremental part is greater than the maintained distance, isValid is set to true (lines 4–6). Then, the TTD $\textit{dist}_{T}$ is computed, which is the distance between the moving object’s target destination $p_{o}^{d}$ and a line segment on the historical trajectory. (line 7) Additionally, the HTD $\textit{dist}_{H}$ is initialized as 00 (line 7). Then, $\textit{dist}_{H}$ is computed (lines 8–14). If the maintained distance is valid, we only update the distance between $p_{o}^{c}$ and $T^{\prime}_{i}$ (Case 2 in Strategy3). Otherwise, $\textit{dist}_{H}$ should be computed by traversing every point $p_{o}^{j}$ in $T_{o}$ (from $p_{o}^{c}$ to previous points) to compute the minimum distance between $p_{o}^{j}$ and $T^{\prime}_{i}$ using Min_Dist function, where the segment pruning (Strategy 1) and the granularity adjustment (Strategy 4) are used during the computation(lines 18–24). During the traverse, the $p_{\textit{max}}$ is updated (line 12) and the $k$-bound pruning (Strategy 2) is used to eliminate irrelevant candidates early (lines 13–14). Then, the OTRD value is computed by summing $\textit{dist}_{H}$ and $\textit{dist}_{T}$ with a trade-off parameter $\alpha$ (line 15). Next, the intermediate buffer is updated (line 16). Finally, the OTRD value and the updated intermediate buffer are returned (line 17).
@@
3:$s_{i}^{z}\leftarrow\textsc{Find\_Pivotal\_Segment}(p_{o}^{c},T_{i})$
4:$p_{i}^{x}\leftarrow\textsc{Find\_Pivotal\_Point}(p_{o}^{c},T_{i})$
@@
9:$\textit{dist}_{T}\leftarrow\textsc{Compute\_Dist\_T}(p_{o}^{d},p_{i}^{x},s_{i}^{z},T_{i})$, $\textit{dist}_{H}\leftarrow 0$
@@
11:    if $p_{o}^{j}\neq p_{o}^{c}$ and $isValid$ then continue $\triangleright$ Strategy 3
12:    $\textit{dist}_{n}\leftarrow$ Min_Dist($j$, 00, $z$, $T$, $\infty$, $g$)
@@
17:$\textit{dist}\leftarrow\alpha\cdot\textit{dist}_{H}+(1-\alpha)\cdot\textit{dist}_{T}$
18:$B_{I}\leftarrow$ Update_B($T,x,z,\textit{dist}_{H},p_{\textit{max}}$)
@@
20:function Min_Dist($j$, $z^{\prime}$, $z$, $T_{i}$, $\textit{dist}_{n}$, $g$)
@@ 5.3. CSTS Algorithm
Algorithm 3 CSTS_Algorithm($o$, $\mathcal{T}$, $r$, $k$, $g$, $SVTI$)
1:Input: A moving object $o$, a historical trajectory set $\mathcal{T}$, a range $r$, an integer $k$, the granularity parameter $g$, segment-based vessel trajectory index $SVTI$.
@@
[…]
@@ 6.1. Overall Experimental Settings
[…]
where $\mathcal{T}_{f}$ is the set of trajectories that cover the $k$-nearest neighbor ($k$NN) points of moving object $o$’s next future point and $\mathcal{T}$ is the query result.
@@
6.2. Evaluation on SVTI
6.2.1. Default
(a) Time vs. Dataset
(b) Size vs. Dataset
Figure 6. Comparison of index construction time and size
6.2.2. Effect of Data Size.
We analyze the effect of varying data sizes on construction time and index size, with results shown in Figures 7(a) and 7(b), respectively. The datasets D-1, D-2, D-3, D-4, and D-5 contain approximately 43, 86, 130, 172, and 216 million points for DK-AIS, and 31, 62, 93, 123, and 154 million points for US-AIS. As the data size increases, the number of segments and corresponding insertions also grow. This leads to a proportional increase in both construction time and index size.
(a) Time vs. $\Lambda$
(b) Size vs. $\Lambda$
Figure 7. Effect of data size
6.2.3. Effect of $l_{\textit{max}}$ and $l_{\text{min}}$
(a) Time vs. $l_{\text{min}}$
(b) Size vs. $l_{\text{min}}$
Figure 8. Effect of $l_{\text{min}}$ ($l_{\textit{max}}=l_{\text{min}}+20$)
(a) Time vs. $l_{\textit{max}}$
(b) Size vs. $l_{\textit{max}}$
Figure 9. Effect of $l_{\textit{max}}$ ($l_{\text{min}}=30$)
@@ 6.3.1. Default
6.3.2. Effect of $l_{q}$
As shown in Figures 11(a) and 11(c), the query times for all methods increase as the query length $l_{q}$ increases. Among the traditional methods, CSTS exhibits the smallest rate of increase because it only evaluates the historical portion of trajectories, avoiding point-by-point checks for the entire trajectory. The query time for O-PQT also increases slightly since it uses the same similarity measure as CSTS, but its reliance on a point-based quad-tree yields higher search costs. H/F-TrajI and H/F-DFT exhibit a more pronounced growth in query time as $l_{q}$ increases, as each point in the historical part of the query trajectory is compared against the entire set of historical trajectories. In contrast, TrajCL-TI and CLEAR-TI show almost no variation in query time varying $l_{q}$, as these methods transform trajectories into fixed-length vectors, making them independent of query length. Figures 11(b) and 11(d) reveal no discernible trend in the hit rate, suggesting that the effect of $l_{q}$ on hit rate is relatively minor.
@@ 6.3.3. Effect of $r$
As shown in Figures 12(a) and 12(c), the query time for all methods increases with the range $r$, as a larger range results in more candidates to evaluate. H/F-TrajI and H/F-DFT experience the fastest growth in query time due to the higher computational cost of similarity computations for each candidate compared to OTRD and the AI-based methods. Among all methods, CSTS remains the most efficient, even as the range increases, owing to its optimized search algorithm. Figures 12(b) and 12(d) show that the hit rate decreases as the range increases. This is because a larger range includes more candidates. Some of these candidates might appear in the final similar trajectory results but not in the $k$NN of $o$’s next position, leading to a reduced hit rate.
@@ 6.3.5. Effect of Data Size
We evaluate CSTS under varying data sizes, as shown in Figure 14. The settings for datasets D-1 through D-5 remain consistent with those described in Section 6.2.2. As the data size increases, both the query time and the hit rate remain relatively stable. This is because the number of candidates remains consistent regardless of the dataset size. Consequently, the query time is primarily determined by the similarity computation rather than the candidate-finding phase, which incurs negligible overhead.
@@ 6.3.6. Effect of $\theta$ and $\alpha$
6.4. Ablation Study

### v2

@@ Abstract.
Tiantian Liu and Hengyu Liu have equal contributions.
@@ 1. Introduction
Automatic Identification System (AIS) data, which provides real-time information on vessel positions, speeds, and directions, offers a rich resource for analyzing vessel movements. With the growing availability of AIS data, understanding vessel trajectories has become essential for navigation safety (Li et al., 2024b) and environmental monitoring (Jin et al., 2021). Trajectory similarity search (Hu et al., 2023) plays a key role in applications such as collision avoidance (Li et al., 2022), route optimization (Xu et al., 2024), and movement pattern analysis (Hu et al., 2024).
@@ 2.1. Trajectory
A segment $s_{i}=\langle p_{i}^{j^{\prime}},\dots,p_{i}^{j}\rangle$ of a trajectory $T_{i}$ is a continuous sub-sequence of $T_{i}$ defined by an index range $j^{\prime},j]$, where $1\leq j^{\prime}\leq j\leq n$.
@@ 2.2. Problem Definition
[…]
@@ 2.3. Framework Overview
Segment-based Vessel Trajectory Index module. Historical trajectories are partitioned into smaller segments using an MBR-optimization method and organized in an index. This index captures connection information, enabling efficient reconstruction of full trajectories when needed and supports spatially optimized search.
Similarity Computation module. It computes the real-time distance between a moving object and historical trajectories using the novel similarity measure OTRD, which consists of the Historical Trajectory Distance (HTD) and the Target-Trajectory Distance (TTD). The OTRD measure prioritizes the future trends of vessel movements and the target destination while also considering the relevance of trajectory segments and the weighted significance of individual points.
@@ 4.2. Object-Trajectory Real-time Distance
[…]
@@ 5.1. Speed-up Strategies
[…]
@@ 5.2. OTRD Computation
8:  if $\textit{dist}_{n}>B_{I}[T_{i}].\textit{dist}^{\prime}_{H}$ then $\textit{isValid}\leftarrow\textit{true}$
@@
11:  if $p_{o}^{j}\neq p_{o}^{c}$ and isValidisValid then continue $\triangleright$ Strategy 3
12:  $\textit{dist}_{n}\leftarrow$ Min_Dist($j$, 0, $z$, $T$, $\infty$, $g$)
@@
[…]
@@
20:function Min_Dist($j$, $^{\prime}, zz$, $T_{i}$, $\textit{dist}_{n}$, $g$)
@@
Algorithm 1 shows the details of the similarity computation. Specifically, we first identify the nearest segment $s_{i}^{z}$ and the nearest point $p_{i}^{x}$ on the historical trajectory $T_{i}$ to the moving object’s current position $p_{o}^{c}$ (lines 1–2). We initialize a boolean flag isValid to check if the pre-computed distance in the intermediate buffer $B_{I}$ is still valid (line 3). A variable $p_{\textit{max}}$ is also initialized to represent the ID of the point $p_{o}^{y}$ in $T_{o}$, where $\theta^{c-y}*\textit{min}\{d(p_{o}^{y},p_{i})|p_{i}\in T^{\prime}_{i}\}=\mathcal{D}_{H}(T_{o},T^{\prime}_{i})$ (line 3). If the pivotal point ID of $T_{i}$ increases and the distance between $p_{o}^{y^{\prime}}$ and the incremental part is greater than the maintained distance, isValid is set to true (lines 4–6). Then, the TTD $\textit{dist}_{T}$ is computed, which is the distance between the moving object’s target destination $p_{o}^{d}$ and a line segment on the historical trajectory. (line 7) Additionally, the HTD $\textit{dist}_{H}$ is initialized as 0 (line 7). Then, $\textit{dist}_{H}$ is computed (lines 8–14). If the maintained distance is valid, we only update the distance between $p_{o}^{c}$ and $T^{\prime}_{i}$ (Case 2 in Strategy3). Otherwise, $\textit{dist}_{H}$ should be computed by traversing every point $p_{o}^{j}$ in $T_{o}$ (from $p_{o}^{c}$ to previous points) to compute the minimum distance between $p_{o}^{j}$ and $T^{\prime}_{i}$ using Min_Dist function, where the segment pruning (Strategy 1) and the granularity adjustment (Strategy 4) are used during the computation(lines 18–24). During the traverse, the $p_{\textit{max}}$ is updated (line 12) and the $k$-bound pruning (Strategy 2) is used to eliminate irrelevant candidates early (lines 13–14). Then, the OTRD value is computed by summing $\textit{dist}_{H}$ and $\textit{dist}_{T}$ with a trade-off parameter $\alpha$ (line 15). Next, the intermediate buffer is updated (line 16). Finally, the OTRD value and the updated intermediate buffer are returned (line 17).
@@ 5.3. CSTS Algorithm
Algorithm 2 CSTS_Algorithm($o$, $\mathcal{T}$, $r$, $k$, $g$, SVTISVTI)
1:Input: A moving object $o$, a historical trajectory set $\mathcal{T}$, a range $r$, an integer $k$, the granularity parameter $g$, segment-based vessel trajectory index SVTISVTI.
@@
11:    $\textit{dist}\leftarrow$ OTRD_Computation($o$, $T_{i}$, $\alpha$, $\theta$, $g$, $\textit{dist}_{k}$, $B_{I}$)
@@ 6.1. Overall Experimental Settings
Metrics. The metrics for each task are listed in Table 3. To evaluate the efficiency of SVTI, we measure its construction time and index size. To evaluate the performance of CSTS, we select ten query instances, execute the selected queries ten times, and report the average. For effectiveness evaluation, we employ the hit rate ($\mathit{Hit}$):
(14) $\mathit{Hit}=\frac{\lvert\mathcal{T}_{f}\cap\mathcal{T}\rvert}{k},$
where $\mathcal{T}_{f}$ is the set of trajectories that cover the $k$-nearest neighbor (kkNN) points of moving object $o$’s next future point and $\mathcal{T}$ is the query result.
@@ 6.2.1. Default
6.2.2. Effect of $l_{q}$
As shown in Figures 7(a) and 7(c), the query times for all methods increase as the query length $l_{q}$ increases. Among the traditional methods, CSTS exhibits the smallest rate of increase because it only evaluates the historical portion of trajectories, avoiding point-by-point checks for the entire trajectory. The query time for O-PQT also increases slightly since it uses the same similarity measure as CSTS, but its reliance on a point-based quad-tree yields higher search costs. H/F-TrajI and H/F-DFT exhibit a more pronounced growth in query time as $l_{q}$ increases, as each point in the historical part of the query trajectory is compared against the entire set of historical trajectories. In contrast, TrajCL-TI and CLEAR-TI show almost no variation in query time varying $l_{q}$, as these methods transform trajectories into fixed-length vectors, making them independent of query length. Figures 7(b) and 7(d) reveal no discernible trend in the hit rate, suggesting that the effect of $l_{q}$ on hit rate is relatively minor.
@@ 6.2.3. Effect of $r$
As shown in Figures 8(a) and 8(c), the query time for all methods increases with the range $r$, as a larger range results in more candidates to evaluate. H/F-TrajI and H/F-DFT experience the fastest growth in query time due to the higher computational cost of similarity computations for each candidate compared to OTRD and the AI-based methods. Among all methods, CSTS remains the most efficient, even as the range increases, owing to its optimized search algorithm. Figures 8(b) and 8(d) show that the hit rate decreases as the range increases. This is because a larger range includes more candidates. Some of these candidates might appear in the final similar trajectory results but not in the kkNN of $o$’s next position, leading to a reduced hit rate.
@@ 6.2.5. Effect of Data Size
We evaluate CSTS under varying data sizes, as shown in Figure 10. The settings for datasets D-1 through D-5 remain consistent with those described in Section B.1.2. As the data size increases, both the query time and the hit rate remain relatively stable. This is because the number of candidates remains consistent regardless of the dataset size. Consequently, the query time is primarily determined by the similarity computation rather than the candidate-finding phase, which incurs negligible overhead.
@@ 6.2.6. Effect of $\theta$ and $\alpha$
7. Related Work
Trajectory similarity measures. Trajectory similarity measures form the foundation of trajectory analysis. Dynamic Time Warping (Yi et al., 1998; Ying et al., 2016) allows for flexible alignment and comparison of time-series data, accommodating variations in speed and temporal shifts. Longest Common Subsequence (Vlachos et al., 2002) offers robustness to noise and spatiotemporal deformations by identifying the longest subsequence shared between trajectories. Additionally, Fréchet distance (Alt and Godau, 1995; De Berg et al., 2013; Agarwal et al., 2014) and edit distance-based measures (Chen and Ng, 2004; Chen et al., 2005; Ranu et al., 2015) provide nuanced assessments of trajectory similarity, considering geometric constraints and structural differences. The Hausdorff distance (Hangouet, 1995; Bai et al., 2011) measures how far the points of one set are from the points of the other set. These measures (Toohey and Duckham, 2015; Magdy et al., 2015) serve as the cornerstone for trajectory analysis. However, these measures fail to consider the significance of individual points, the trends within historical trajectories, and the future movement patterns of objects.
Learning-based trajectory similarity computation. Recent advancements in trajectory similarity search have explored learning-based approaches (Li et al., 2018; Fu and Lee, 2020; Liu et al., 2020; Deng et al., 2022; Fang et al., 2022; Mao et al., 2022; Li et al., 2024a) to improve the accuracy and efficiency of similarity computations. These methods often leverage deep learning models to automatically learn representations of trajectories from raw spatiotemporal data. However, learning-based trajectory similarity computation has several limitations that make it less suitable for real-time, continuous vessel trajectory search. First, these methods often rely on down-sampling or distortion of trajectories, which leads to the identification of trajectories as similar, even when they lack true alignment with the original paths. This compromises their ability to accurately capture the underlying similarity between the original trajectories. Second, these methods typically focus on computing the similarity between two given trajectories rather than provide efficiently similarity search method on a large trajectory database. Third, they are not well-suited for real-time processing. The continuous updates in vessel trajectories would necessitate frequent model re-training or recalibration, imposing substantial computational overhead and significantly degrading performance. Finally, learning-based models are generally designed for offline batch processing, making them unsuitable for dynamic and real-time maritime navigation.
Similar trajectory search. Studies on similar trajectory search (Ding et al., 2008; Xie et al., 2017; Wang et al., 2018; Li et al., 2020; He et al., 2022; Luo and Zeng, 2023) and similarity join (Bakalov et al., 2005b, a; Ta et al., 2017; Shang et al., 2017) have proposed various techniques to identify similar movement patterns. Some approaches (Hwang et al., 2006, 2005; Xie et al., 2017; Yuan and Li, 2019; Han et al., 2021; He et al., 2022; Fang et al., 2022) are widely used in domains like vehicle tracking and urban traffic monitoring, particularly in road networks, where vehicle movements are constrained by structured paths, making it easier to organize and index trajectories based on fixed road segments, but not suitable for vessel trajectories. Furthermore, they mainly designed for existing trajectory similarity methods to compare entire trajectories, which is not suitable for real-time similar vessel trajectory search. These methods are less effective for real-time applications, where the trajectory evolves over time and must be analyzed incrementally. As a result, current trajectory similarity search methods are not well-suited to solve the problem of real-time, continuous vessel trajectory search.
8. Conclusion and Future Work
Acknowledgements.
We acknowledge the support of the EU’s Horizon Europe research program under grant agreement No. 101070279 (MobiSpaces).
Appendix A Algorithm of Segment Partitioning
Algorithm 3 outlines the segment partitioning process in detail. Initially, a segment set $\mathcal{S}$ is initialized (line 1). For each trajectory $T_{i}$, the algorithm identifies the optimal segmentation (lines 2–16). Specifically, a list is created for $T_{i}$’s segment set, where the $j$th item represents the current optimal segment set up to point $p^{j}$ (line 3). Additionally, another list is initialized to store the total MBR area, where the $j$th item represents the minimum total MBR area up to point $p^{j}$ (line 4). For each point $p^{j}$ in $T_{i}$, the minimum total MBR area up to $p^{j}$ is computed (lines 5–15). If $j$ is less than $l_{\textit{min}}$, the area is set to 0; otherwise, it is initialized to $\infty$ (lines 6–7). For each possible starting point $p^{j-l+1}$, where $l_{\textit{min}}\leq l\leq l_{\textit{max}}$, the MBR area of the segment $\langle p^{j-l+1},...,p^{j}\rangle$ is computed (lines 8–10). If the total MBR area is smaller than the current value, both the area and the segment set are updated (lines 11–15). Once all points in $T_{i}$ have been processed, the final segment set for $T_{i}$ is added to the overall segment set $\mathcal{S}$ (line 16). Ultimately, the algorithm returns the overall segment set $\mathcal{S}$ (line 17).
2:Output: A segment set $\mathcal{S}$.
3:Initialize a segment set $\mathcal{S}$ $\leftarrow$ $\emptyset$
4:for $T_{i}=\langle p_{i}^{1},p_{i}^{2},...,p_{i}^{n}\rangle$ in $\mathcal{T}$ do
5:  Initialize a list for $T_{i}$’s segment set $S_{i}$ $\leftarrow$ $\emptyset$
6:  Initialize a list for the total MBR area $F$ $\leftarrow$ $\emptyset$
7:  for $j$ in $[1,n]$ do
8:   if $j\leq l_{\textit{min}}$ then $F[j]$ $\leftarrow$ 0
9:   else $F[j]$ $\leftarrow$ $\infty$
10:   for $l$ in $l_{\textit{min}},l_{\textit{max}}]$ do
11:     if $j-l\geq 0$ then
12:      $m_{l}$ $\leftarrow$ $\mathcal{M}_{area}$($p_{i}^{j-l+1}$, $p_{i}^{j}$)
13:      if jlj-l is 0 then $m$ $\leftarrow$ $m_{l}$
14:      else $m$ $\leftarrow$ $F[j-l]+m_{l}$
15:      if $F[j]>m$ then
16:        $F[j]$ $\leftarrow$ $m$
17:        $S_{i}[j]$ $\leftarrow$ $S_{i}[j-l]$ $\cup$ $\langle p_{i}^{j-l+1},...,p_{i}^{j}\rangle$
18:  $\mathcal{S}\leftarrow\mathcal{S}\cup S_{i}[n]$
19:return $\mathcal{S}$
Appendix B Evaluation
B.1. Evaluation on SVTI
B.1.1. Default
(a) Time vs. Dataset
(b) Size vs. Dataset
Figure 12. Comparison of index construction time and size
B.1.2. Effect of Data Size.
We analyze the effect of varying data sizes on construction time and index size, with results shown in Figures 13(a) and 13(b), respectively. The datasets D-1, D-2, D-3, D-4, and D-5 contain approximately 43, 86, 130, 172, and 216 million points for DK-AIS, and 31, 62, 93, 123, and 154 million points for US-AIS. As the data size increases, the number of segments and corresponding insertions also grow. This leads to a proportional increase in both construction time and index size.
(a) Time vs. $\Lambda$
(b) Size vs. $\Lambda$
Figure 13. Effect of data size
B.1.3. Effect of $l_{\textit{max}}$ and $l_{\text{min}}$
(a) Time vs. $l_{\text{min}}$
(b) Size vs. $l_{\text{min}}$
Figure 14. Effect of $l_{\text{min}}$ ($l_{\textit{max}}=l_{\text{min}}+20$)
(a) Time vs. $l_{\textit{max}}$
(b) Size vs. $l_{\textit{max}}$
Figure 15. Effect of $l_{\textit{max}}$ ($l_{\text{min}}=30$)
B.2. Ablation Study